OpenStudy (aravindg):

TOUGH ONE!!! The value of p for which the roots of the quadratic equation 4x^2-20px+(25p^2+15p-66)=0 are less than 2 ,lies in interval .......................

OpenStudy (mathteacher1729):

Oh there's some fun algebra here... :) Ultimately this is a quadratic equation: \[ax^2 + bx + c = 0\] Where \[a = 4 \qquad b = -20p \qquad c = 25p^2+15p-66\] Now you have to use the quadratic formula to find out where the roots are less than 2. Let me try to cook up a geogebra file to visualize this.

OpenStudy (aravindg):

show me

hero (hero):

c = OMG

OpenStudy (aravindg):

wat?

hero (hero):

nvm

OpenStudy (mathteacher1729):

I've attached a geogebra file which illustrates this. It doesn't give an exact solution of course, but hopefully it lends some intuition to the problem. :)

OpenStudy (aravindg):

its not working plz give algebraic answer

hero (hero):

Wait, let me....

OpenStudy (mathteacher1729):

You need geogebra to open it. http://www.geogebra.org/ and click installers. It's well worth it. :)

OpenStudy (mathteacher1729):

Arivand, did my hint help? How far have you gotten with this problem so far? I will better be able to help this way.

OpenStudy (aravindg):

plz give algebraic answer

hero (hero):

{x = 3 / 4, x = (-3) / 4}

hero (hero):

I used geogebra and found that solution. mathteacher, you could have used CAS on Geogebra to find it

OpenStudy (mathteacher1729):

Arivind, I am happy to help you work through the problem, but I don't just give solutions without any student input. There is no learning that way. :)

OpenStudy (aravindg):

plz give steps

OpenStudy (anonymous):

4x^2-20px+(25p^2+15p-66)=0 here , we use quadratic equation where , a =4 b= -20p and, c = (25p^2+15p-66)

OpenStudy (aravindg):

thn............

OpenStudy (aravindg):

how?

hero (hero):

You should be happy you have the solutions. You can use the quadratic formula instead

OpenStudy (anonymous):

now sum of roots is \[\alpha +\beta=-(-20p)/4\]

OpenStudy (aravindg):

so.............

OpenStudy (anonymous):

\[\alpha +\beta=5p\]

OpenStudy (anonymous):

\[\alpha \times \beta= c/a\]

OpenStudy (aravindg):

................ lead me to answer